Circle Circumfe Circle Diameter r is radius Finish A sprinter runs the curve of this 200 m in 10.34 s. Assume he ran in a lane which makes a semicircle (r 36.3 m) for the first part of the race. At 3.44 s into the race his speed is 4.0 m/s. At 8.0 s into the race his speed is 9.6 m/s. His speed after the curve was 12.8 m/s. What was the sprinters angular velocity at 3.44 s? Answer in deg/s
Answer #1

From the relation between the linear and angular quantities we know that

V=r*W

V is linear velocity

r is radius of the curvaacur

W is angular velocity

Substituting the values

W= V/r

W= 4/36.3 rad/s

W= 0.1102 rad/s

So the angular velocity at t=3.44s is W= 0.1102rad/s